The equation 𝑧^2 = −5 + 12𝑖 has two (complex) solutions. Note that
𝑧 = sqrt (−5 + 12𝑖)is not a particularly useful way to write a solution, since it does not tell us the real and imaginary parts of the solution. We are actually trying to find solutions in the form 𝑧 = 𝑥 + 𝑖𝑦. Show that 𝑧 = 2 + 3𝑖 is such a solution and find the second solution. I know how to prove that z= 2 +3i is a solution, but I just don't know how to find the other solution? Please explain!

Answer by rothauserc(4718)

𝑧^2 = −5 + 12𝑖
:
To solve this problem, we use the polar form of a complex number and De Moivre's Theorem
:
The polar form of a complex number(a+bi, where a,b are real numbers) is r(cos theta + i * sin theta) where r > or = to zero, r is the distance formula = square root(a^2+b^2) and tan theta = b/a
:
write 𝑧^2 = −5 + 12𝑖 in polar form
:
r = square root((-5)^2 + 12^2) = square root(25 + 144) = 13
tan theta = 12/-5 = -2.4 and theta is -67.38
Note that 180 - 67.38 is 112.62
:
then z^2 = 13(cos (112.62) + i * sin(112.62))
:
using De Moivre's Theorem, we know that z^2 has two roots
:
let n be a positive number(2 for this problem) and let w(k) be the kth root, then
:
w(k) = r^(1/2)[ cos((112.62 + 2kpi)/2) + ( i * sin((112.62 + 2kpi/2)))
where k = 0, 1 an pi = 180 degrees
:
w(0) = 13(1/2) * (cos(112.62 / 2) + (i * sin(112.62 / 2))
w(0) = 2 + 3i
:
w(1) = 13(1/2) * (cos((112.62+360) / 2) + (i * sin((112.62+360) / 2))
w(1) = -2 -3i
:
*****************************
the other solution is -2 -3i
*****************************
:
: